A friction less pulley of radius 0.4m a rope over there and a lift is moving upward with acc 0.6m/S2 for 10 seconds then how much round will pulley take.

Grade:

3 Answers

Eshan
5 years ago
Dear student,

Distance travelled by lift in given time=Length of rope that made rounds around the pulley

=s=\dfrac{1}{2}at^2=30m
Circumference of the pulley=2\pi R
Hence the number of rounds pulley makes=\dfrac{30}{2\pi R}=11.9\approx 12
Aman Mittal
5 years ago
Sir why we not take g'=g+0.6 as it is moving with upward acceleration as you take acceleration 0.6m/S2 only
Eshan
5 years ago
Dear Aman,

The effective acceleration due to gravity as experienced by a person standing inside the lift will definitely beg+a. But the number of rounds that the pulley will make has nothing to do with the acceleration of gravity as felt by man inside lift. All we want to know is the real distance travelled by the lift so that we can find out the length of wrapped wire on pulley, hence it is the net acceleration of the lift that we are required to find out.

Hope it helps. Regards.